Solve the system
3x+4y=23
2x-5y=23

Answer by checkley71(8403)

3x+4y=23 multiply by 5:
15x+20y=115 2x-5y=23 multiply by 4
8x-20y=92 now add
----------------------------------
23x=207
x=207/23
x=9 answer.
3*9+4y=23
27+4y=23
4y=23-27
y=-4/4
y=-1 answer.
proof
2*9-5-1=23
18+5=23
23=23